I am supposed to write a quadratic equation in the variable x as having the given numbers as a solution.Type the equation in standard form ax^2 + bx + c =0
Solution: 6 is the only solution

Answer by ankor@dixie-net.com(12684)

I am supposed to write a quadratic equation in the variable x as having the given numbers as a solution.
Type the equation in standard form ax^2 + bx + c = 0
Solution: 6 is the only solution
:
FOIL: (x-6)(x-6) = x^2 - 6x - 6x + 36 = x^2 - 12x + 36 = 0